OpenStudy (anonymous):

Simplify each expression (square root sign)125

OpenStudy (anonymous):

\[\sqrt{125} = \sqrt{5 \times 5 \times 5} = \sqrt{5^{2} \times 5} = \sqrt{5^{2}} \times \sqrt{5} = 5\sqrt{5}\]